Circular No. 6522

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996by in UGC 3379 Weidong Li, Qiran Qiao, Yulei Qiu, and Jingyao Hu, Beijing Astro-

nomical Observatory (BAO), on behalf of the BAO Supernova Survey, report their discovery of a supernova located 7′′.2 west and 27′′.9 south of the nucleus of UGC 3379 (α = 5^{h}58^{m}26^{s}, δ = +68^{o}27′.6, equinox 2000.0). Unfiltered CCD images taken on Dec. 14 and 15 with the BAO 0.60-m reflector show the object at mag 16.5 and 16.1, respectively. CCD images of the same field taken on Dec. 7 and 8 (limiting mag ∼ 19.0) and on Dec. 11 (limiting mag ∼ 18.0) show no star at the position of SN 1996by. A low-resolution, high-quality spectrogram (0.48 nm/pixel, range 400–800 nm), taken with the BAO 2.16-m telescope by W. Li, J. Wei, L. Cao, and Q. Qiao on Dec. 15.75 UT shows that this is a type-Ia supernova prior to optical maximum. There is strong Si II 635.5-nm absorption at 607 nm. Strong lines of S II, Fe III, and Ca II are also present, with minima measured at 574.0, 538.6, 481.4, 423.5, and 393.6 nm. The photospheric expansion velocity, measured from the minimum of the Si II doublet, is ∼ 14 600 km/s.

2S 0114+650 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association, for the RXTE ASM team at GSFC and the Massachusetts Institute of Technology; and J. P. Finley, Purdue University, report: “RXTE ASM observations of 2S 0114+650 obtained between 1996 Jan. 5 and Dec. 5 show the presence of a modulation at a period of 2.7388 ± 0.0006 hr and a peak-to-peak amplitude of ∼ 60 percent. We identify this with the 2.78-hr (± 0.01 hr) period reported by Finley et al. (1992, A.Ap. 262, L25), showing that this is a persistent coherent property of the source. In addition, the ASM data indicate a modulation at a period of 11.7 ± 0.2 days (amplitude ∼ 60 percent), consistent with the 11.6-day optical period (Crampton et al. 1985, Ap.J. 299, 839). These results can be interpreted in terms of a neutron star rotating with a 2.7-hr period in an 11.6-day eccentric orbit around its early-type companion.”

COMET 81P/WILD 2 Total (visual) magnitude estimates: Nov. 6.80 UT, 14.4 (A. Nakamura,

Kuma, Ehime, Japan, 0.60-m reflector + unfiltered CCD); 22.14, 13.5: (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); Dec. 4.33, 12.5 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1996 December 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green